The retina of a human eye can detect light when radiant energy incident on it is at least 4.0X10-17 J

2. The retina of a human eye can detect light when radiant energy incident on it is at least 4.0X10-17 J. For light of 600-nm wavelength, how many photons does this correspond to?
Answer #1
Photon energy = h*c/(wavelength)
= 6.62*10^-34 J*s * (3*10^8 m/s)/600*10^-9 m = 3.3*10^-19 J

4*10^-17 J corresponds to 120 photons of that color (orange) light
answered by: Megan